# Exploring Selective Layer Fine-Tuning in Federated Learning

### v2

@@ Abstract
Federated learning (FL) has emerged as a promising paradigm for fine-tuning foundation models using distributed data in a privacy-preserving manner. Under limited computational resources, clients often find it more practical to fine-tune a selected subset of layers, rather than the entire model, based on their task-specific data. In this study, we provide a thorough theoretical exploration of selective layer fine-tuning in FL, emphasizing a flexible approach that allows the clients to adjust their selected layers according to their local data and resources. We theoretically demonstrate that the layer selection strategy has a significant impact on model convergence in two critical aspects: the importance of selected layers and the heterogeneous choices across clients. Drawing from these insights, we further propose a strategic layer selection method that utilizes local gradients and regulates layer selections across clients. The extensive experiments on both image and text datasets demonstrate the effectiveness of the proposed strategy compared with several baselines, highlighting its advances in identifying critical layers that adapt to the client heterogeneity and training dynamics in FL.
@@ 1 Introduction
This study focuses on selective model fine-tuning as it is particularly well-suited to address the inherent heterogeneity in FL, i.e., the data heterogeneity and device heterogeneity (Yang et al., 2021; Chai et al., 2019; Li et al., 2022). Specifically, clients involved in FL have non-independent and identically distributed (non-IID) data distribution and different system resources, leading to the need to customize fine-tuning strategies to such discrepancies. For example, clients with limited computation resources may opt to update only a fraction of the model, while those with sufficient resources and high-quality data prefer fine-tuning a large portion of the model to enhance performance. Selective model fine-tuning enables clients to adjust the chosen part of the model to be updated based on their own capabilities, providing a flexible and advanced solution to mitigating sub-optimal issues induced by the heterogeneity in FL.
The exploration of selective model fine-tuning within the context of FL, is still in its early stages. Previous studies (Shen et al., 2021; Xu et al., 2021; Lee et al., 2022; Dun et al., 2022) have concentrated on designing static strategies for subnetwork selection to improve model fine-tuning performance, without adequately considering heterogeneity among clients. To fulfill this gap, in this paper, we provide a comprehensive theoretical analysis on selective model fine-tuning in FL, focusing on a general scenario where clients are allowed to choose different layers for local training and vary their choices across different training epochs, called selective layer fine-tuning. Specifically, we formulate the optimization objective of selective layer fine-tuning in FL, and provide insights on effectively determining critical layers to achieve model convergence. Building on these insights, we further propose a novel layer selection strategy that leverages local gradients and the regulation of unified selections.
@@
  • We conduct experiments to compare the proposed layer selection strategy with baseline methods on both image and text datasets. The experimental results demonstrate the superior performance of the proposed strategy in achieving better model performance, highlighting that the proposed strategy can find critical layers for fine-tuning while considering the client heterogeneity and training dynamics in FL.
@@ 2 Related Works
Recent studies have extended the selective fine-tuning techniques to FL scenarios (Nguyen et al., 2022a; Chen et al., 2022; Hilmkil et al., 2021; Zhang et al., 2022b). Specifically, researchers (Lee et al., 2023; Dun et al., 2022) investigate layer-wise network decomposition to achieve selective model fine-tuning on clients. However, these works fail to offer methodologies for adaptive and dynamic layer selection that take into account the heterogeneous characteristics of clients. In addition, personalized FL algorithms (Pillutla et al., 2022; Chen et al., 2023b) propose to train different subnetworks on clients towards better local models. Different from previous studies, we focus on providing an in-depth understanding of selective layer fine-tuning in FL, considering the heterogeneity from the perspective of client resources and local data distributions.
@@ 4 Which Layers Should Be Selected For Fine-Tuning?
These strategies for layer selection are proposed based on the insights drawn from experts’ experience, serving as special instantiations of the selected layer sets $\mathcal{L}_{i}^{t}$. It is worth noting that these experience-driven strategies might not consistently yield optimal results in various FL applications, particularly considering the client heterogeneity. This leads to an essential question: How to effectively determine the task-specified layer selection strategy among a large search space of possible options? In the rest of this section, we provide a theoretical analysis to answer this question.
@@ 4.1 Theoretical Analysis
$\vspace{-5pt}\mathcal{E}_{t}\leq 2\underbrace{\Bigg{[}\Bigg{\|}\sum_{l\notin\mathcal{L}_{t}}\nabla f(\theta^{t})\Bigg{\|}^{2}\Bigg{]}}_{\mathcal{E}_{t,1}}+2\underbrace{\sum_{l\in\mathcal{L}_{t}}\chi_{\mathbf{w}_{t,l}\|\mathbf{\alpha}}\kappa_{l}^{2}}_{\mathcal{E}_{t,2}},$ (14)
@@ 4.2 Layer Selection Strategy
4.3 Discussions on Computational and Communication costs
In this section, we provide discussions on the computational cost and communication costs, considering a case where each client selects $R$ layers to fine-tune a model with $L$ layers.
Computational cost   Since all methods require the same forward operations, we focus on comparing the computational cost of backward operations in different methods. For simplicity, we assume each layer requires $b$ FLOPs of backward operations. The average computational cost of the proposed layer selection method is calculated as:
$\text{Cost}_{\text{sel}}=\underbrace{b(L-1)}_{\text{Select}}+\underbrace{bR\tau}_{\text{Fine-tune}}=b(R\tau+L-1),$ (16)
where $\tau$ represents the local training steps. For comparison, fully fine-tuning a model requires the computational cost of:
$\text{Cost}_{\text{full}}=bL\tau=\frac{L\tau}{R\tau+L-1}\text{Cost}_{\text{sel}}.$ (17)
Thus, the proposed method takes a much lower computational cost than the full model fine-tuning, and the cost reduction is proportional to the number of layers and local training steps.
Meanwhile, the layer selection step in the proposed method introduces slightly additional costs of $\frac{L-R}{\tau L}\text{Cost}_{\text{full}}$, which can be further reduced by evaluating the model on a smaller volume of data or making the selection decision at a lower frequency.
Communication time   The communication time is determined by the transmitted bits of different methods. The proposed method only transmits the selected layers, which is much less than uploading the full model. For example, assuming that different layers have the same number of parameters, the communication cost of the proposed method is $\frac{R}{L}$ of full model fine-tuning.
To summarize, the computational cost and communication cost (time) of the proposed method are much lower than the full model fine-tuning. In the following section, the empirical observations also verify the benefits of the proposed selective layer fine-tuning method.
@@ 5.3 Further Discussions
Comparisons regarding computational and communication costs   We compare the computational cost (in TFLOPs) and communication cost (in transmitted MBits) of the proposed method with the full model fine-tuning when adopting the CLIP model on the CIFAR-10. For the proposed method, we consider fine-tuning only one layer, as it is sufficient to achieve comparable accuracy with full model fine-tuning according to Table 1. The results in Table 3 evidence a substantial decrease in both computational and communication requirements when utilizing the proposed method. Besides, the layer selection strategy takes as low as 3.4% of the computational cost. Therefore, the aforementioned empirical observations demonstrate that the proposed method significantly reduces both the computational and communication costs.
Table 3: Comparisons of computational and communication costs when fine-tuning the CLIP model on the CIFAR-10 ($R=1$). The numbers in brackets represent the cost of the layer selection strategy.
Computational cost (TFLOPs) Ratio Transmission (MBits) Ratio
Full Model Fine-tuning 8.47 100% 2,811 100%
Proposed Method 2.24 (1.51) 26% (17%) 234 8.33%
Proposed (Sel. Period=2) 1.46 (0.75) 17% (9.5%) 234 8.33%
Proposed (Sel. Batch=1) 0.99 (0.30) 12% (3.4%) 234 8.33%

### v3

@@ Abstract
Federated learning (FL) has emerged as a promising paradigm for fine-tuning foundation models using distributed data in a privacy-preserving manner. Under limited computational resources, clients often find it more practical to fine-tune a selected subset of layers, rather than the entire model, based on their task-specific data. In this study, we provide a thorough theoretical exploration of selective layer fine-tuning in FL, emphasizing a flexible approach that allows the clients to adjust their selected layers according to their local data and resources. We theoretically demonstrate that the layer selection strategy has a significant impact on model convergence in two critical aspects: the importance of selected layers and the heterogeneous choices across clients. Drawing from these insights, we further propose a strategic layer selection method that utilizes local gradients and regulates layer selections across clients. Extensive experiments on both image and text datasets demonstrate the effectiveness of the proposed strategy compared with several baselines, highlighting its advances in identifying critical layers that adapt to the client heterogeneity and training dynamics in FL.
@@ 1 Introduction
This study focuses on selective model fine-tuning as it is particularly well-suited to address the inherent heterogeneity in FL, i.e., the data heterogeneity and device heterogeneity (Yang et al., 2021; Chai et al., 2019; Li et al., 2022). Specifically, clients involved in FL have non-independent and identically distributed (non-IID) data and different system resources, leading to the need to customize fine-tuning strategies to such discrepancies. For example, clients with limited computation resources may opt to update only a fraction of the model, while those with sufficient resources and high-quality data prefer fine-tuning a large portion of the model to enhance performance. Selective model fine-tuning enables clients to adjust the chosen part of the model to be updated based on their own capabilities, providing a flexible and advanced solution to mitigating sub-optimal issues induced by the heterogeneity in FL.
The exploration of selective model fine-tuning within the context of FL, is still in its early stage. Previous studies (Shen et al., 2021; Xu et al., 2021; Lee et al., 2022; Dun et al., 2022) have concentrated on designing static strategies for subnetwork selection to improve model fine-tuning performance, without adequately considering heterogeneity among clients. To fulfill this gap, in this paper, we provide a comprehensive theoretical analysis on selective model fine-tuning in FL, focusing on a general scenario where clients are allowed to choose different layers for local training and vary their choices across different training epochs, called selective layer fine-tuning. Specifically, we formulate the optimization objective of selective layer fine-tuning in FL, and provide insights on effectively determining critical layers to achieve model convergence. Building on these insights, we further propose a novel layer selection strategy that leverages local gradients and the regulation of unified selections.
@@
  • We conduct experiments to compare the proposed layer selection strategy with baseline methods on both image and text datasets. Experimental results demonstrate the superior performance of the proposed strategy in achieving better model performance, highlighting that the proposed strategy can find critical layers for fine-tuning while considering the client heterogeneity and training dynamics in FL [^1] The source codes are available at https://github.com/hiyuchang/fed_sel_tune..
@@ 2 Related Works
Recent studies have extended the selective fine-tuning techniques to FL scenarios (Nguyen et al., 2022a; Chen et al., 2022; Hilmkil et al., 2021; Zhang et al., 2022b). Specifically, researchers (Lee et al., 2023; Dun et al., 2022) investigate layer-wise network decomposition to achieve selective model fine-tuning on clients. However, these works fail to offer methodologies for adaptive and dynamic layer selection that takes into account the heterogeneous characteristics of clients. In addition, personalized FL algorithms (Pillutla et al., 2022; Chen et al., 2023b) propose to train different subnetworks on clients towards better local models. Different from previous studies, we focus on providing an in-depth understanding of selective layer fine-tuning in FL, considering the heterogeneity from the perspective of client resources and local data distributions.
@@ 4 Which Layers Should Be Selected For Fine-Tuning?
These strategies for layer selection are proposed based on the insights drawn from experts’ experience, serving as special instantiations of the selected layer sets $\mathcal{L}_{i}^{t}$. It is worth noting that these experience-driven strategies might not consistently yield optimal results in various FL applications, particularly considering client heterogeneity. This leads to an essential question: How to effectively determine the task-specified layer selection strategy among a large search space of possible options? In the rest of this section, we provide a theoretical analysis to answer this question.
@@ 4.1 Theoretical Analysis
$\vspace{-5pt}\mathcal{E}_{t}\leq 2\underbrace{\Bigg{[}\Bigg{\|}\sum_{l\notin\mathcal{L}_{t}}\nabla_{l}f(\theta^{t})\Bigg{\|}^{2}\Bigg{]}}_{\mathcal{E}_{t,1}}+2\underbrace{\sum_{l\in\mathcal{L}_{t}}\chi_{\mathbf{w}_{t,l}\|\mathbf{\alpha}}\kappa_{l}^{2}}_{\mathcal{E}_{t,2}},$ (14)